OpenStudy (shawnbrown1209):

Barney wants to rent a tent. He has to pay a fixed base cost plus a daily rate for renting the tent. The table shows the amount of money, y, in dollars, that Barney has to pay for renting the tent for x days: Number of days(x) Rent (dollars)(y) (x) (y) 0 3 1 8 2 13 3 18 4 23 Which equation best shows the relationship between x and y? (A)y = 5x + 3 (B) y = x + 8 (C) y = 5x + 8 (D) y = x + 3

OpenStudy (phi):

do you know how to find the slope if you have two points ?

OpenStudy (shawnbrown1209):

@phi I tried to do the work but I don't really understand

OpenStudy (shawnbrown1209):

Y1-Y2 --------- X1-X2

OpenStudy (phi):

ok, pick out two points (x,y) pairs from the table

OpenStudy (shawnbrown1209):

You're giving me the choice to pick any number from the table? @phi

OpenStudy (phi):

yes, you can use any pair. But I would use the first two (smaller numbers!)

OpenStudy (shawnbrown1209):

Okay. Then, yeah, I choose the first two.

OpenStudy (phi):

find the "change in y" and the "change in x"

OpenStudy (shawnbrown1209):

tbh I honestly don't understand any of this (pls don't judge me Im so so bright in algebra)

OpenStudy (shawnbrown1209):

not so**

OpenStudy (phi):

write down the two pairs: (0,3), (1, 8) find the difference (i.e. subtract) the y values. what do you get ?

OpenStudy (shawnbrown1209):

(0,3) - (1,8) = 1.5

OpenStudy (phi):

when you have time, see https://www.khanacademy.org/math/algebra-basics/core-algebra-graphing-lines-slope/core-algebra-slope/v/slope-of-a-line-2 you first do 8-3 = 5 and then you do 1-0 = 1 then 5/1 which is change in y divided by change in x 5/1 is just 5

OpenStudy (shawnbrown1209):

dang you're judging me hard sorry

OpenStudy (phi):

so the equation is y= 5x + b to "find b" I would put in x=0 and y= 3 (i.e. use the first pair) 3 = 5*0 + b 5 times 0 is 0 3= 0 + b adding 0 does not change b, so you get 3= b